OpenStudy (anonymous):

a doubly charged helium atom is accelerated by the voltage V=500V in the right bound direction and then it enters the region with the electric field E=200V/cm directed to the left. By what distance L will the atom penetrate the region of the field?

OpenStudy (anonymous):

Studying for a physics 2 final and having a hard time relating everything. I'm trying to find the charge through E=V/q then once i have the charge I can use E=kq/d^2 to find my distance or length . I converted the v/cm to v/m, But I'm getting a really crazy length. 3000 meters. I know that's probably not correct

OpenStudy (anonymous):

doubly charged helium +2 = 2(-e) - electron charge. so Energy = 2*electron charge *volt. = 0.5 m v^2 --->1 Thus you have the initial Velocity. and when Helium enters the electric field a force due to this E will affect him in opposite direction to its motion so F = q E and q = 2(-e). thus a = f/m. so you will get the initial velocity from 1 and the final velocity will be 0 and v^2 = 2 a x so you can solve it for x :D

OpenStudy (anonymous):

For anyone else that happens to be following along, did we end up with 2.5x10^-12m?